OpenStudy (adrianna.....ramirez):

x + y = 3k x - y = k The solution of the system shown is _____.

jhonyy9 (jhonyy9):

x+y=3k x-y=k --------- 2x 0 = 4k x= 4k/2 = 2k x=2k x+y=3k 2k +y=3k y = 3k-2k = k y=k x=2k y=k

OpenStudy (mathmale):

jhonny is using the "elimination by addition and subtraction method" here, first to find x, and then to find k.